# Bayesian Self-Distillation for Image Classification

### v1

@@ Abstract
Supervised training of deep neural networks for classification typically relies on hard targets, which promote overconfidence and can limit calibration, generalization, and robustness. Self-distillation methods aim to mitigate this by leveraging inter-class and sample-specific information present in the model’s own predictions, but often remain dependent on hard targets, reducing their effectiveness. With this in mind, we propose Bayesian Self-Distillation (BSD), a principled method for constructing sample-specific target distributions via Bayesian inference using the model’s own predictions. Unlike existing approaches, BSD does not rely on hard targets after initialization. BSD consistently yields higher test accuracy (e.g. +1.4% for ResNet-50 on CIFAR-100) and significantly lower Expected Calibration Error (ECE) (-40% ResNet-50, CIFAR-100) than existing architecture-preserving self-distillation methods for a range of deep architectures and datasets. Additional benefits include improved robustness against data corruptions, perturbations, and label noise. When combined with a contrastive loss, BSD achieves state-of-the-art robustness under label noise for single-stage, single-network methods. The source code is available at https://github.com/antonadelow/BSD.
@@ 1 Introduction
Despite the widespread use of deep neural networks for classification tasks, the implications of using hard targets for loss computation have received relatively little attention. Intuitively, classes can exhibit varying degrees of similarity, and individual samples may exhibit different levels of resemblance to both their assigned and other classes. This nuanced, sample-specific information – often referred to as dark knowledge (hinton2014dark) – is not taken into account when relying solely on hard targets. For instance, misclassifying an image of a dog as a teapot is fundamentally different from misclassifying it as a cat. Nevertheless, deep neural networks are typically trained with hard targets, which fail to capture inter-class relationships or sample-specific ambiguities, contributing to poor calibration (guo2017calibration).
Self-distillation, an efficient variant of knowledge distillation (hinton2015distilling), leverages dark knowledge by using a model’s own predictions as soft targets during training. In contrast, conventional knowledge distillation provides soft targets to a model by leveraging another, often larger, network. Existing self-distillation methods, however, exhibit varying limitations. Some modify the network architecture with intermediate classifiers (zhang2019your; zhang2022self), increasing parameter count and computational cost. Others impose a consistency loss between current predictions and those from the last epoch or mini-batch (kim2021self; shen2022self). Yet, epoch-wise construction introduces variance under augmentations, while targets from the last mini-batch may be of limited effectiveness, as they originate from nearly identical model states.
@@
Fundamentally, existing self-distillation frameworks constrain their predictions by incorporating a loss term derived from the hard targets (e.g. furlanello2018born; kim2021self; shen2022self). This reliance interferes with the goal of learning richer and more calibrated predictions, and leaves the model sensitive to label noise. In the overparameterized regime, sensititivity to label noise can exacerbate double descent (nakkiran2021deep), a phenomenon where the test error initially decreases, then increases, before decreasing again as model capacity or training time increase. We hypothesize that a self-distillation method decoupled from the hard targets during training could provide smoother and more robust probability estimates, thereby mitigating these effects and providing more predictable training dynamics.
To address these limitations, we propose Bayesian Self-Distillation (BSD), an efficient method for leveraging dark knowledge by constructing rich, sample-specific target distributions in a single training run. BSD treats a model’s own predictions as evidence for Bayesian inference, without explicitly relying on the hard targets after initialization. A diagram depicting BSD is included in Figure 1.
@@
  • We propose Bayesian Self-Distillation (BSD), a lightweight and principled method for self-distillation that operates independently of hard targets after initialization.
  • We provide the first, to our knowledge, theoretical formalization of dark knowledge and use it to show that BSD captures rich sample-specific information.
  • We demonstrate through extensive experiments that BSD consistently outperforms both existing architecture-preserving self-distillation methods and conventional knowledge distillation in generalization and calibration for a variety of datasets and network architectures.
  • We show that BSD provides robustness against data corruptions, perturbations, and label noise. Notably, it mitigates epoch-wise double descent under label noise and achieves state-of-the-art performance for single-stage, single-network methods when combined with a contrastive loss.
@@ 2 Related work
Self-Distillation (SD) originates from Knowledge Distillation (KD) (hinton2015distilling), a technique introduced for model compression using a teacher–student framework. Later work demonstrated that identical (Born-Again) networks can be trained sequentially with KD to improve generalization (furlanello2018born), thereby giving rise to self-distillation.
@@
SD methods differ primarily in how they construct the teacher signal. One family of approaches relies on architectural structure, such as introducing auxiliary classifiers or branches to provide internal supervision (zhang2019your; zhang2022self; lan2018knowledge). Another category leverages temporal knowledge transfer, using model snapshots from earlier epochs (yang2019snapshot), moving averages of past predictions (Temporal Ensembling, TE) (laine2016temporal), or moving averages of model parameters (mean teacher) (tarvainen2017mean) over epochs. A third set of methods leverages consistency, encouraging alignment between samples of the same class (yun2020regularizing) or between consecutive predictions of the same sample (kim2021self; shen2022self). This category includes Progressive Self-Knowledge Distillation (PS-KD) (kim2021self), which uses the model’s prediction from the last epoch as a soft target, and Self-Distillation from the Last mini-Batch (DLB) (shen2022self), which constructs overlapping mini-batches and leverages the prediction from the previous mini-batch. Despite their differences, these strategies remain anchored to the original hard labels, which can be particularly problematic in settings where labels are noisy.
@@
Label noise remains an obstacle in deep learning, contributing to overfitting and the undesirable training dynamics. Proposed solutions range from regularization techniques such as label smoothing (szegedy2016rethinking) and Mixup (zhang2017mixup), to robust loss functions like Symmetric Cross Entropy (wang2019symmetric). More sophisticated methods include sample selection via optimal transport feng2023ot; chang2023csot, label correction by interpolation between the prediction and label (xu2025revisiting), or even complex multi-stage (liu2023chimera) or multi-network training pipelines (zhang2024bpt). These methods, however, often rely on heuristics rather than reformulating the learning objective to explicitly model uncertainty. In contrast, probabilistic modeling provides a principled way to estimate uncertainty and improves interpretability.
@@
Bayesian methods in deep learning have gained attention for their ability to model predictive uncertainty. Methods like Variational inference (blundell2015weight), Monte Carlo dropout (gal2016dropout), and deep ensembles (lakshminarayanan2017simple) can be used to quantify uncertainty, improve calibration and increase robustness to out-of-distribution data. Recently, Bayesian optimization of hyperparameters has been paired with born-again networks to improve generalization over generations of networks. Related work has also shown that KD itself can serve as a mechanism to transfer better calibration from a teacher to a student (hebbalaguppe2024calibration).
Our proposed method, Bayesian Self-Distillation (BSD), draws inspiration from these domains. Like SD, it softens target distributions, but views the problem from a Bayesian perspective where model predictions serve as evidence for updating target distributions. Note that BSD is distinct from both Bayesian neural networks and Bayesian optimization, where the former regards distributions over model weights and the latter is a method for optimizing black-box functions, whereas BSD places priors over target distributions and refines them over epochs. This allows targets to evolve independently without the continual influence of hard targets, providing richer and more flexible self-supervision that inherently offers robustness to label noise.
@@ 3 Methodology
As discussed in Section 1, a network trained exclusively on hard targets is hardly encouraged to learn meaningful inter-class relationships or account for sample-specific ambiguities. To address this, we frame the training process as Recursive Bayesian Estimation, with the goal of approximating the true latent target distributions. Rather than viewing the network as a static generator of fixed predictions, we treat it as a noisy sensor observing the external dataset. By recursively integrating these noisy signals, we aim to construct stable target distributions that capture both class relationships and sample-specific information, thereby improving generalization and reducing overconfidence. The proposed method, Bayesian Self-Distillation (BSD), is summarized in Algorithm 1.
@@ 3.1 Notation
Consider the supervised classification problem with a dataset $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}^{0}_{i})\}_{i=1}^{n},$ samples $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and one-hot targets $\mathbf{y}^{0}_{i}\in\Delta_{k}$. Let $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$ denote a neural network with parameters $\bm{\theta}$ and a softmax activation as its last layer, $\mathcal{L}$ the loss function, and let $\hat{\mathbf{y}}_{i}^{t}$ denote the prediction the model outputs for sample $i$ at epoch $t$.
3.2 Bayesian Self-Distillation
Training a deep neural network is a stochastic process where randomness enters through stochastic optimization (e.g., SGD), data augmentation, and regularization mechanisms. SGD dynamics approximate a posterior distribution over model parameters (mandt2017stochastic), while stochastic regularization (e.g. dropout) allows the network’s output to be viewed as a draw from a predictive distribution (gal2016dropout). Consequently, the model’s prediction for a sample $\mathbf{x}_{i}$ at any training step can be seen as a random sample from an implicit predictive distribution.
From this perspective, we treat the network as a noisy sensor measuring the true latent class distribution $\mathbf{y}_{i}$ of the external input $\mathbf{x}_{i}$. This interpretation is similar to established frameworks in that it views neural network training as Recursive Bayesian Estimation (singhal1988training). At epoch $t$, the network provides a prediction $\hat{\mathbf{y}}_{i}^{t}$, serving as a noisy measurement of the true state $\mathbf{y}_{i}$. Although the sensor evolves during training making the measurement noise non-stationary, this dynamic is consistent with adaptive filtering and probabilistic self-training frameworks (e.g., Expectation-Maximization). Importantly, this framing allows for model predictions to be viewed as external and independent evidence. The predictions $\hat{\mathbf{y}}_{i}^{t}$ constitute measurements as read by the sensor $f$, conditioned on the independent input $\mathbf{x}_{i}$. To formalize this, we introduce a latent class variable $z_{i}\in\{1,\ldots,k\}$ for each sample $\mathbf{x}_{i}$ and model its distribution as categorical,
@@ 3.2 Bayesian Self-Distillation
To express prior beliefs about $\mathbf{y}_{i}$, we use a Dirichlet distribution $\mathbf{y}_{i}\sim\text{Dir}(\bm{\alpha}_{i})$ due to its conjugacy and representation of accumulated evidence as independent components. Specifically, a Dirichlet distribution can be generated by normalizing a set of independent Gamma variables $v_{i,j}\sim\text{Gamma}(\alpha_{i,j},1)$ such that $y_{i,j}=v_{i,j}/\sum_{k}v_{i,k}$. This allows us to interpret the prior as maintaining independent evidence counters for each class. When the input $\mathbf{x}_{i}$ triggers a non-zero prediction for a secondary class (e.g., “cat” features in a “dog” image), it acts as an independent sensor update for that specific class. Assuming that the model learns to extract relevant feature representations during training, the predictions will capture semantically meaningful knowledge rather than model hallucinations. We encode prior belief in the class corresponding to the label by letting
@@
which generalizes the categorical likelihood to fractional evidence and is conjugate to the prior (bishop2006pattern). This likelihood treats the prediction as partial evidence, where higher predicted probabilities correspond to stronger observations.
Formally, after observing a prediction $\hat{\mathbf{y}}_{i}^{t}$ at epoch $t$, the posterior distribution is
@@
which accumulates the noisy measurements into sample-specific distributions over the class probabilities $\mathbf{y}$. To predict the distribution over the labels, we use the posterior predictive distribution. However, because the quality of measurements is expected to improve during training (as the model learns), the measurement noise is non-stationary. Standard Bayesian updating would weight early, noisy observations equally to later, more accurate ones. To address this, we adopt a discounted Bayesian model (west2006chapter7), effectively forgetting old evidence to adapt to the improving network. For a discounting factor $\gamma\in[0,1]$, we have $\bm{\alpha}_{i}^{t}=\gamma\bm{\alpha}^{t-1}_{i}+\hat{\mathbf{y}}_{i}^{t}$, yielding the update rules
$\mathbf{y}^{t}_{i}=\frac{\gamma A_{i}^{t-1}}{\gamma A_{i}^{t-1}+1}\mathbf{y}_{i}^{t-1}+\left(1-\frac{\gamma A_{i}^{t-1}}{\gamma A_{i}^{t-1}+1}\right)\hat{\mathbf{y}}_{i}^{t},\quad A_{i}^{t}=\gamma A_{i}^{t-1}+1,$ (5)
@@
Intuitively, the Dirichlet parameters accumulate the model’s belief about how much each class is supported by the input $\mathbf{x}_{i}$. Discounting ensures that more recent predictions are weighted more heavily than earlier, likely worse, ones. Early in training, when predictions are more likely to have high variance, the prior dominates. As training progresses, predictions are expected to become more consistent, providing stronger and more reliable evidence.
Algorithm 1 Bayesian Self-Distillation (BSD)

### v2

# Deep Probabilistic Supervision for Image Classification

@@ Abstract
Supervised training of deep neural networks for classification typically relies on hard targets, which promote overconfidence and can limit calibration, generalization, and robustness. Self-distillation methods aim to mitigate this by leveraging inter-class and sample-specific information present in the model’s own predictions, but often remain dependent on hard targets without explicitly modeling predictive uncertainty. With this in mind, we propose Deep Probabilistic Supervision (DPS), a principled learning framework constructing sample-specific target distributions via statistical inference on the model’s own predictions, remaining independent of hard targets after initialization. We show that DPS consistently yields higher test accuracy (e.g., +2.0% for DenseNet-264 on ImageNet) and significantly lower Expected Calibration Error (ECE) (-40% ResNet-50, CIFAR-100) than existing self-distillation methods. When combined with a contrastive loss, DPS achieves state-of-the-art robustness under label noise. The source code is available at https://github.com/antonadelow/DPS.
@@ 1 Introduction
Despite the widespread use of deep neural networks for classification tasks, the implications of using hard targets for loss computation have received relatively little attention. Intuitively, classes can exhibit varying degrees of similarity, and individual samples may enjoy different levels of resemblance to both their assigned and other classes. This nuanced, sample-specific information – often referred to as dark knowledge (Hinton, 2014) – is discarded when relying solely on hard targets. For instance, misclassifying an image of a dog as a teapot is inherently different from misclassifying it as a cat. Nevertheless, deep networks are typically trained with hard targets, which fail to capture inter-class relationships or sample-specific ambiguities, contributing to poor calibration (Guo et al., 2017).
Self-distillation, an efficient variant of knowledge distillation (Hinton et al., 2015), leverages dark knowledge by combining a model’s own predictions with the hard targets during training. In contrast, conventional knowledge distillation provides soft targets to a model by leveraging another, often larger, network. Despite their benefits, existing self-distillation methods exhibit varying limitations. Some modify the network architecture with intermediate classifiers (Zhang et al., 2019, 2021), increasing parameter count and computational cost. Others impose a consistency loss between current predictions and those from the last epoch or mini-batch (Kim et al., 2021; Shen et al., 2022). Yet, epoch-wise construction introduces variance under augmentations, while targets from the last mini-batch may be of limited effectiveness, as they originate from nearly identical model states.
@@
Fundamentally, existing self-distillation frameworks do not model predictive uncertainty explicitly, which may affect calibration negatively, while also constraining their predictions by incorporating a loss term derived from the hard targets (e.g. Furlanello et al. (2018); Kim et al. (2021); Shen et al. (2022)). This reliance interferes with the goal of learning more calibrated and nuanced predictions, leaving the model sensitive to label noise. In the overparameterized regime, sensitivity to label noise can exacerbate double descent (Nakkiran et al., 2021), a phenomenon where the test error initially decreases, then increases, before decreasing again as model capacity or training time varies. In this work, we hypothesize that a self-distillation method decoupled from hard targets could provide smoother and more robust probability estimates, thereby mitigating these effects and providing more predictable training dynamics.
To address these limitations, we propose Deep Probabilistic Supervision (DPS), an efficient learning framework for leveraging dark knowledge by constructing nuanced, sample-specific target distributions in a single training run by modeling predictive uncertainty explicitly. DPS treats a model’s own predictions as evidence for Bayesian inference, without relying on hard targets after initialization. A diagram depicting DPS is included in Figure 1.
@@
  • We propose Deep Probabilistic Supervision (DPS), a lightweight and principled framework for constructing nuanced target distributions during training that operates independently of hard targets after initialization.
  • We demonstrate through extensive experiments that DPS consistently outperforms both existing self-distillation methods and conventional knowledge distillation in generalization and calibration.
  • We show that DPS provides robustness against data corruptions, perturbations, and label noise. Notably, DPS mitigates epoch-wise double descent under label noise and achieves state-of-the-art performance when combined with a contrastive loss.
  • We provide the first, to our knowledge, theoretical formalization of dark knowledge and use it to show that DPS captures nuanced, sample-specific, information.
@@ 2 Related work
The notion of Self-Distillation (SD) originates from Knowledge Distillation (KD) (Hinton et al., 2015), a technique introduced for model compression using a teacher–student framework. Later work demonstrated that identical (Born-Again) networks can be trained sequentially with KD to improve generalization (Furlanello et al., 2018), thereby giving rise to self-distillation.
@@
SD methods differ primarily in how they construct the teacher signal. One family of approaches relies on architectural structure, such as introducing auxiliary classifiers or branches to provide internal supervision (Zhang et al., 2019, 2021; Zhu et al., 2018), while another leverages temporal knowledge transfer, using model snapshots from earlier epochs (Yang et al., 2019), or moving averages of either past predictions (Temporal Ensembling, TE) (Laine and Aila, 2016) or model parameters (Mean Teacher) (Tarvainen and Valpola, 2017). A third category enforces consistency, either between same-class samples (Yun et al., 2020) or consecutive predictions, including methods such as Progressive Self-Knowledge Distillation (PS-KD) (Kim et al., 2021) and Self-Distillation from the Last mini-Batch (DLB) (Shen et al., 2022) that leverage predictions from the previous epoch or mini-batch, respectively. Despite their differences, these strategies remain anchored to the original hard labels, which can be particularly problematic in settings where labels are noisy.
@@
Label noise remains an obstacle in deep learning, contributing to overfitting and undesirable training dynamics. Proposed solutions range from regularization techniques such as label smoothing (Szegedy et al., 2016) and Mixup (Zhang et al., 2017), to robust loss functions like Symmetric Cross Entropy (Wang et al., 2019). More sophisticated methods include sample selection via optimal transport Feng et al. (2023); Chang et al. (2023), label correction by interpolation between the prediction and label (Xu et al., 2025), or complex multi-stage (Liu et al., 2023) or multi-network training pipelines (Zhang et al., 2024), but often rely on heuristics rather than reformulating the learning objective to explicitly model uncertainty. Probabilistic modeling provides a principled way to estimate uncertainty, improving interpretability.
@@
Bayesian methods in deep learning have gained attention for their ability to model predictive uncertainty. Methods like Variational inference (Blundell et al., 2015), Monte Carlo dropout (Gal and Ghahramani, 2016), and deep ensembles (Lakshminarayanan et al., 2017) are used to quantify uncertainty, improve calibration or increase robustness to out-of-distribution data. Related work has also shown that KD itself can serve as a mechanism to transfer better calibration from teacher to student (Hebbalaguppe, 2024).
Our proposed framework, Deep Probabilistic Supervision (DPS), is inspired by these domains. Like SD, DPS constructs softened target distributions, but models target construction probabilistically by leveraging model predictions as evidence. Distinct from other Bayesian methods that model distributions over weights or predictions, DPS models distributions over targets. DPS allows targets to evolve independently of the original labels, offering flexible self-supervision that inherently models uncertainty and increases robustness against label noise.
@@ 3 Methodology
As discussed in Section 1, a network trained exclusively on hard targets is not incentivized to learn meaningful inter-class relationships or account for sample-specific ambiguities. To address this, and explicitly model uncertainty in target distributions, we frame the training process as Recursive Bayesian Estimation (Singhal and Wu, 1988). Rather than viewing the network as a static generator of fixed predictions, we treat it as a sensor observing the external dataset under the stochastic dynamics of training, and aim to construct stable target distributions by recursively integrating predictions for improved generalization and reduced overconfidence. The proposed method, Deep Probabilistic Supervision (DPS), is summarized in Algorithm 1.
@@ 3.1 Notation
Consider the supervised classification problem with a dataset $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}^{0}_{i})\}_{i=1}^{n}$ consisting of samples $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and one-hot targets $\mathbf{y}^{0}_{i}\in\Delta_{k}$. Let $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$ denote a neural network with parameters $\bm{\theta}$ and a softmax activation as its last layer, let $\ell$ be the loss function, and let $\hat{\mathbf{y}}_{i}^{t}$ denote the prediction the model outputs for sample $i$ at epoch $t$.
3.2 Deep Probabilistic Supervision
Training deep networks is a stochastic process where randomness enters through stochastic optimization (e.g., SGD), data augmentation, and regularization mechanisms. SGD dynamics approximate a posterior distribution over model parameters (Mandt et al., 2017), while predictions under stochastic regularization have been suggested to represent draws from an implicit predictive distribution (Gal and Ghahramani, 2016). Consequently, we consider the model’s prediction of a sample $\mathbf{x}_{i}$ as a random sample from a predictive distribution.
From this perspective, we treat the network as a sensor measuring the true class distribution $\mathbf{y}_{i}$ of the external input $\mathbf{x}_{i}$, an interpretation similar to established frameworks in that it views neural network training as Recursive Bayesian Estimation (Singhal and Wu, 1988). Although the sensor evolves during training, making the measurement noise non-stationary, this dynamic is consistent with adaptive filtering and probabilistic self-training frameworks (e.g., Expectation-Maximization). Importantly, model predictions $\hat{\mathbf{y}}_{i}^{t}$ can be viewed as independent evidence, since they constitute measurements as read by the sensor $f$, conditioned on external input $\mathbf{x}_{i}$. To formalize this, we introduce a latent class variable $z_{i}\in\{1,\ldots,k\}$ for each sample $\mathbf{x}_{i}$ and model its distribution as categorical,
@@ 3.2 Deep Probabilistic Supervision
To express prior beliefs about $\mathbf{y}_{i}$, we use a Dirichlet distribution $\mathbf{y}_{i}\sim\text{Dir}(\bm{\alpha}_{i})$ for its conjugacy and its construction from normalized independent Gamma variables (if $v_{i,j}\sim\text{Gamma}(\alpha_{i,j},\lambda)$ then Dirichlet variates are obtained by the closure $y_{i,j}=v_{i,j}/\sum_{k}v_{i,k})$. This formulation allows us to interpret the accumulated evidence for the classes as independent, so that when the input $\mathbf{x}_{i}$ triggers a non-zero prediction for a secondary class (e.g., “cat” features in a “dog” image), the component for that class receives an independent sensor update. Assuming that the model learns to extract relevant features, the predictions capture semantic knowledge rather than hallucinations. We encode prior belief in the labeled class by letting
@@
which generalizes the categorical likelihood to fractional evidence, and is conjugate to the prior (Bishop and Nasrabadi, 2006). This likelihood treats the prediction as partial evidence, where higher predicted probabilities correspond to stronger observations.
Formally, the posterior distribution is
@@
accumulating noisy measurements into sample-specific distributions over class probabilities $\mathbf{y}$. To predict the distribution over the labels, we use the posterior predictive distribution. However, because the quality of measurements is expected to improve during training (as the model learns), the measurement noise is non-stationary. Standard Bayesian updating would weight early, noisier observations equally to later, more accurate ones. To address this, we adopt a discounted Bayesian model (West and Harrison, 2006), effectively forgetting old evidence to adapt to the improving network. For a discounting factor $\gamma\in[0,1]$, we have $\bm{\alpha}_{i}^{t}=\gamma\bm{\alpha}^{t-1}_{i}+\hat{\mathbf{y}}_{i}^{t}$, yielding the update rules
$\mathbf{y}^{t}_{i}=\frac{\gamma A_{i}^{t-1}}{\gamma A_{i}^{t-1}+1}\mathbf{y}_{i}^{t-1}+\left(1-\frac{\gamma A_{i}^{t-1}}{\gamma A_{i}^{t-1}+1}\right)\hat{\mathbf{y}}_{i}^{t},\quad A_{i}^{t}=\gamma A_{i}^{t-1}+1,\quad$ (5)
@@
From the perspective of Bayesian filtering, discounting the concentration parameters $\bm{\alpha}_{i}^{t-1}$ constitutes the prediction step, increasing the variance of the distribution, while adding the new evidence $\hat{\mathbf{y}}_{i}^{t}$ forms the update step Särkkä and Svensson (2023).
Intuitively, the Dirichlet parameters accumulate the model’s belief about how much each class is supported by the input $\mathbf{x}_{i}$. Discounting ensures that more recent predictions are weighted more heavily than earlier, likely worse, ones. The prior dominates early in training when variance is high, but predictions provide stronger evidence as training progresses.
3.3 Relationship between DPS and other methods
Related methods can be interpreted as special cases of DPS. Conventional training corresponds to taking $c\to\infty$, which fixes the target distribution and prevents Bayesian updating. In this limit, the framework reduces to standard distribution matching, as KL-divergence and cross-entropy are equivalent under one-hot targets.
Algorithm 1 Deep Probabilistic Supervision (DPS)